# Non-stationary Spatio-Temporal Modeling Using the Stochastic Advection-Diffusion Equation

### v1

@@ Abstract
We construct flexible spatio-temporal models through stochastic partial differential equations (SPDEs) where both diffusion and advection can be spatially varying. Computations are done through a Gaussian Markov random field approximation of the solution of the SPDE, which is constructed through a finite volume method. The new flexible non-separable model is compared to a flexible separable model both for reconstruction and forecasting and evaluated in terms of root mean square errors and continuous rank probability scores. A simulation study demonstrates that the non-separable model performs better when the data is simulated with non-separable effects such as diffusion and advection. Further, we estimate surrogate models for emulating the output of a ocean model in Trondheimsfjorden, Norway, and simulate observations of autonoumous underwater vehicles. The results show that the flexible non-separable model outperforms the flexible separable model for real-time prediction of unobserved locations.
@@ 1 Introduction
Spatio-temporal data consists of observations that have associated both spatial locations and time points. For example, temperature or precipitation measured at different weather stations, or the concentration of pollutants in the ocean measured from ships. Developing new statistical models for analyzing such data is an active area of research (Porcu et al.,, 2021). The overarching goal is first to accurately capture the complex interactions between space and time in natural phenomena while also enabling efficient computation for handling large datasets. Then, to leverage the estimated interactions to predict unobserved locations or to forecast.
@@
A common approach to model spatio-temporal data is through Gaussian random fields (GRFs); see, for example, Cressie and Wikle, (2011). GRFs are fully characterized by a mean function and a covariance function, which means that the key questions are how to model the mean structure and how to model the covariance structure? However, the covariance function must be positive semi-definite, and this complicates the construction of valid models with realistic features.
@@
A simple way to guarantee a valid spatio-temporal covariance function is to take the product of a spatial covariance function and a temporal covariance function. This is called a separable covariance function but requires a very specific and simple interaction between space and time, and natural processes such as advection and diffusion are non-separable. Typically, one would construct separable covariance functions using stationary and isotropic covariance functions such as the Matérn family of covariance functions, which is popular due to the interpretable parameters and the flexibility in controlling the smoothness, range, and variance. The specification of covariance functions for highly dynamic and non-stationary processes is known to be challenging (Cressie and Wikle,, 2011; Simpson et al.,, 2012), and assumptions such as isotropy, stationarity, and separability are often made to improve computational tractability even though they are physically unrealistic.
@@
Spatio-temporal datasets can be large, and direct computations are limited by a complexity of (${\cal O}(n^{3})$), where $n$ is the number of observations. This is often called the ”big $n$ problem” (Banerjee et al.,, 2003). Many approaches have been suggested to handle large datasets (Heaton et al.,, 2019). Some include include low-rank matrices (Cressie and Johannesson,, 2008; Banerjee et al.,, 2008; Wikle,, 2010), approximate likelihoods (Vecchia,, 1988; Eidsvik et al.,, 2014; Katzfuss and Guinness,, 2021), and Gaussian Markov random fields (GMRFs) (Rue and Held,, 2005). However, an issue with GMRFs is that they are defined through Markov properties and conditional distributions, and are difficult to construct, in general. Lindgren et al., (2011) proposed to overcome this issue by defining models using stochastic partial differential equations (SPDEs) instead of covariance functions, and then constructing GMRFs that approximate the solutions of the SPDEs.
@@
SPDEs are a powerful modeling tool for constructing non-separable spatio-temporal GRFs as they can be related to partial differential equations that describe physical laws. For instance, the stochastic heat or diffusion equation (Heine,, 1955; Jones and Zhang,, 1997; Lindgren et al.,, 2023), the stochastic wave equation (Carrizo Vergara et al.,, 2022), and the stochastic advection-diffusion equation (Liu et al.,, 2022; Clarotto et al.,, 2023). In the context of the SPDE approach, a common way to construct separable covariance structures have been to use an autoregressive process of order 1 (AR(1) process), where the temporal innovations are spatial GRFs (Rodríguez-Iturbe and Mejía,, 1974; Cameletti et al.,, 2013). Spatial non-stationarity can be introduced in the innovations using SPDEs with spatially varying coefficients (Fuglstad et al., 2015a, ; Fuglstad et al., 2015b, ; Hildeman et al.,, 2021; Berild and Fuglstad,, 2023). This gives a separable spatio-temporal covariance structure with spatial non-stationarity (Fuglstad and Castruccio,, 2020).
@@
A promising approach to construct a class of non-separable covariance models with spatial non-stationarity is to start with the stochastic advection-diffusion equation. Previous work has approximated solutions through spectral representations (Sigrist et al.,, 2014; Liu et al.,, 2022), or numerical methods such as finite element methods (FEMs) (Clarotto et al.,, 2023; Lindgren et al.,, 2023). An issue with FEM is that it can be unstable in an advection-dominated setting (Clarotto et al.,, 2023). However, a more natural choice in a conservation law setting is to use the finite volume method (FVM), as it is based on the flow of matter between cells in the discretization, and therefore, the FVM is often closer to the physics of the problem (LeVeque,, 2002). The FVM has been used in Fuglstad et al., 2015a ; Fuglstad et al., 2015b and Berild and Fuglstad, (2023) in a purely spatial setting, but the FVM has not been used in spatio-temporal statistics even though it is prevalent in the field of computational fluid dynamics.
@@
The behavior of the ocean is complex, and we need high-resolution deterministic numerical models that are based on the Navier-Stokes equations to describe it well. Such deterministic models are hard to update in real-time due to their computational complexity (Liu et al.,, 2022). In contrast, statistical models are highly effective for space-time interpolation and prediction, making them suitable for real-time applications. The challenge lies in sufficiently describing the complex spatio-temporal behavior. We can take advantage of the complex numerical models to estimate complex statistical surrogate models using the abundance of numerically simulated data. Many authors have employed such an approach to estimate models used in ocean monitoring with AUVs but with some limitations and simplifications. Ge et al., (2023) estimated a purely spatial model through covariance functions for ocean salinity field in 3D and used an adaptive sampling strategy to choose locations for the autonomous underwater vehicles (AUVs) to sample. Foss et al., (2022) used the output from a numerical model to estimate an advection-diffusion model with covariate-based advection and constant isotropic diffusion for monitoring mine tailings in the ocean. Berild et al., (2024) estimated a non-stationary and anisotropic spatial model through the SPDE approach for a 3D ocean salinity field.
In Section 2, we outline the theoretical foundation for spatio-temporal modeling using SPDEs. Then in Section 3 we introduce the proposed method for modeling spatio-temporal data and elaborate on the parameterization of the stochastic advection-diffusion equation. In Section 4, we demonstrate the proposed method on a synthetic dataset with a complex spatio-temporal structure and compare its predictive performance to other spatio-temporal models. In Section 5, we apply our method to emulate a numerical ocean model, simulate an ocean monitoring scenario with AUVs, and compare the accuracy of the proposed model to a non-stationary separable spatio-temporal model. Finally, Section 6 presents a discussion of the results and our conclusions. Some technical details such as the derivation of the numerical solution of the stochastic advection-diffusion equation are deferred to the appendices. The computer code used in this work is available at https://github.com/berild/spdepy.
@@ 2.1 Non-stationary spatial modelling
where $\kappa(\cdot)$ is a positive function, $\nabla$ is the gradient, and $\mathbf{H}(\cdot)$ is a differentiable spatially varying symmetric positive definite $2\times 2$ matrix, and $\mathcal{W}(\cdot)$ is Gaussian white noise. See Bolin and Kirchner, (2020) for more details on the technical requirements on $\kappa(\cdot)$ and $\mathbf{H}(\cdot)$. We apply zero-flow Neumann boundary conditions,
@@
In the case that $\kappa(\cdot)$ and $\mathbf{H}(\cdot)$ are spatially varying, there is no obvious way to avoid boundary effects since there is no unique way to extend $\kappa(\cdot)$ and $\mathbf{H}(\cdot)$ to a larger domain than the area of interest. In practice, one can use a larger domain than the area of interest, and then estimate $\kappa(\cdot)$ and $\mathbf{H}(\cdot)$ also outside the area of interest (Fuglstad et al., 2015b, ). The anisotropic Laplacian in Equation (1) induces local anisotropy in the correlation at every location, and the combination of $\kappa(\cdot)$ and $\mathbf{H}(\cdot)$ controls the spatially varying spatial dependence and spatially varying marginal variance. See the discussion in Fuglstad et al., 2015b .
In this paper, we will follow Fuglstad et al., 2015a and parametrize $2\times 2$ symmetric positive matrices through
@@ 2.2 Separable spatio-temporal modelling
$\frac{\partial}{\partial t}u(\bm{s},t)+cu(\bm{s},t)=\sqrt{2c}\frac{\partial}{\partial t}B(\bm{s},t),\quad\bm{s}\in\mathcal{D},\quad t\in(0,\infty)$ (5)
@@ 2.3 The stochastic advection-diffusion equation
$\left(\mathbf{H}_{\mathrm{E}}(\bm{s})\nabla u(\bm{s},t)+\bm{\omega}_{\mathrm{E}}(\bm{s})u(\bm{s},t)\right)\cdot{\bm{n}}(\bm{s})=0,\quad\bm{s}\in\partial\mathcal{D},\quad t\in(0,\infty),$ (7)
@@ 3.2 Parametrizing the advection-diffusion-reaction model
where $B_{x,i}(x)$ and $B_{y,j}(y)$ are the 2nd-order B-spline basis functions in the $x$ and $y$ directions, respectively. A visualization of the B-spline basis functions is shown in Figure 1. This is similar to the approach taken by Fuglstad et al., 2015b and Berild and Fuglstad, (2023) on the Whittle-Matérn SPDE. This gives in total 90 coefficients for the basis functions, and thus 91 parameters including $\log(\tau)$. In general, one could use an arbitrary number of B-spline basis functions in each dimension at the cost of increased computation time. We refer readers to Berild and Fuglstad, (2023, Appendix A.3) for a detailed description of the construction of 1D clamped B-splines.
@@
We denote the parameter vector as $\bm{\theta}$, and the vector contains $\log(\tau)$ and all coefficients of basis functions. When $\bm{\theta}$ is known, all functions are known, and the stochastic advection-diffusion-reaction is fully specified. Before we state the full hierarchical model and discuss computations, we discretize the stochastic advection-diffusion-reaction model in space and time.
@@ 3.3 Discretizing the advection-diffusion-reaction model
We solve advection-diffusion-reaction SPDE given in Equation (6) numerically by first discretizing in time by backward Euler, and then in space using a finite volume method.
@@
The spatial domain ${\cal D}$ is subdivided into $K$ regular rectangular grid cells, each represented by a centroid $\bm{s}_{k}$. We assume a piece-wise constant solution with constant value within each grid cell, as is common in standard FVM. A visualization of the spatial discretization is presented in Figure S1. We refer to Appendix A. for technical details and give an overall description of the results below.
@@
${\bm{w}}^{n}=\begin{bmatrix}u^{n}(\bm{s}_{1}),\ldots,u^{n}(\bm{s}_{K})\end{bmatrix}^{\mathrm{T}},\quad n=0,\ldots,T-1.$
@@
$\left(\mathbf{D}_{\mathrm{V}}+\Delta t\left(\mathbf{D}_{\mathrm{V}}\mathbf{D}_{\kappa^{2}}-\mathbf{A}_{\mathbf{H}}+\mathbf{A}_{\bm{\omega}}\right)\right){\bm{w}}^{n+1}=\mathbf{D}_{\mathrm{V}}{\bm{w}}^{n}+\tau\sqrt{\Delta t}\mathbf{D}_{\mathrm{V}}\bm{\phi}^{n},$ (12)
@@
  • $\mathbf{D}_{\mathrm{V}}=V\mathbf{I}_{K}$: a diagonal matrix containing the areas of the cells;
@@
  • $\mathbf{A}_{\bm{\omega}}$: uses an up-wind scheme for advection and details are provided in Appendix A..

### v2

@@ Abstract
We construct flexible spatio-temporal models through stochastic partial differential equations (SPDEs) where both diffusion and advection can be spatially varying. Computations are done through a Gaussian Markov random field approximation of the solution of the SPDE, which is constructed through a finite volume method. The new flexible non-separable model is compared to a flexible separable model both for reconstruction and forecasting, and evaluated in terms of root mean square errors and continuous rank probability scores. A simulation study demonstrates that the non-separable model performs better when the data is simulated from a non-separable model with diffusion and advection. Further, we estimate surrogate models for emulating the output of a ocean model in Trondheimsfjorden, Norway, and simulate observations of autonomous underwater vehicles. The results show that the flexible non-separable model outperforms the flexible separable model for real-time prediction of unobserved locations.
@@ 1 Introduction
Spatio-temporal data consist of observations that have associated both spatial locations and time points. For example, temperature or precipitation measured at different weather stations, or the concentration of pollutants in the ocean measured from ships. Developing new statistical models for analyzing such data is an active area of research (Porcu et al.,, 2021). The overarching goal is first to accurately capture the complex interactions between space and time in natural phenomena while also enabling efficient computation for handling large datasets; then, to leverage the estimated interactions to predict unobserved locations or to forecast.
@@
A common approach to model spatio-temporal data is through Gaussian random fields (GRFs); see, for example, Cressie and Wikle, (2011). GRFs are fully characterized by a mean function and a covariance function, which means that the key challenges are to model the mean structure and to model the covariance structure. However, the covariance function must be positive semi-definite, and this complicates the construction of valid models with realistic features.
@@
A simple way to guarantee a valid spatio-temporal covariance function is to take the product of a spatial covariance function and a temporal covariance function. This is called a separable covariance function, but requires a very specific and simple interaction between space and time, and the natural processes advection and diffusion typically leads to non-separability. Typically, one would construct separable covariance functions using stationary and isotropic covariance functions such as the Matérn family of covariance functions, which is popular due to the interpretable parameters and the flexibility in controlling the smoothness, range, and variance. The specification of covariance functions for highly dynamic and non-stationary processes is known to be challenging (Cressie and Wikle,, 2011; Simpson et al.,, 2012), and assumptions such as isotropy, stationarity, and separability are often made to improve computational tractability even though they are physically unrealistic.
@@
Spatio-temporal datasets can be large, and direct computations are limited by a complexity of ${\cal O}(n^{3})$, where $n$ is the number of observations. This is often called the ”big $n$ problem” (Banerjee et al.,, 2003). Many approaches have been suggested to handle large datasets (Heaton et al.,, 2019). Some include include low-rank matrices (Cressie and Johannesson,, 2008; Banerjee et al.,, 2008; Wikle,, 2010), approximate likelihoods (Vecchia,, 1988; Eidsvik et al.,, 2014; Katzfuss and Guinness,, 2021), and Gaussian Markov random fields (GMRFs) (Rue and Held,, 2005). However, an issue with GMRFs is that they are defined through Markov properties and conditional distributions, and are difficult to construct, in general. Lindgren et al., (2011) proposed to overcome this issue by defining models using stochastic partial differential equations (SPDEs) instead of covariance functions, and then constructing GMRFs that approximate the solutions of the SPDEs.
@@
SPDEs can be a powerful modeling tool for constructing non-separable spatio-temporal GRFs as they can be related to partial differential equations that describe physical laws. For instance, the stochastic heat or diffusion equation (Heine,, 1955; Jones and Zhang,, 1997; Lindgren et al.,, 2023), the stochastic wave equation (Carrizo Vergara et al.,, 2022), and the stochastic advection-diffusion equation (Liu et al.,, 2022; Clarotto et al.,, 2024). In the context of the SPDE approach, a common way to construct separable covariance structures have been to use an autoregressive process of order 1 (AR(1) process), where the temporal innovations are spatial GRFs (Rodríguez-Iturbe and Mejía,, 1974; Cameletti et al.,, 2013). Spatial non-stationarity can be introduced in the innovations using SPDEs with spatially varying coefficients (Fuglstad et al., 2015a, ; Fuglstad et al., 2015b, ; Hildeman et al.,, 2021; Berild and Fuglstad,, 2023). This gives a separable spatio-temporal covariance structure with spatial non-stationarity (Fuglstad and Castruccio,, 2020).
@@
A promising approach to construct a class of non-separable covariance models with spatial non-stationarity is to start with the stochastic advection-diffusion equation. Previous work has approximated solutions through spectral representations (Sigrist et al.,, 2014; Liu et al.,, 2022), or numerical methods such as finite element methods (FEMs) (Clarotto et al.,, 2024; Lindgren et al.,, 2023). An issue with FEM is that it can be unstable in an advection-dominated setting (Clarotto et al.,, 2024). However, a more natural choice in a conservation law setting is to use the finite volume method (FVM) (Eymard et al.,, 2000), as it is based on the flow of matter between cells in the discretization, and therefore, the FVM is often closer to the physics of the problem (LeVeque,, 2002). The FVM has been used in Fuglstad et al., 2015a ; Fuglstad et al., 2015b and Berild and Fuglstad, (2023) in a purely spatial setting, but it has not been used in spatio-temporal statistics even though it is prevalent in the field of computational fluid dynamics.
In this paper, we build on existing work on spatial non-stationarity (Fuglstad et al., 2015b, ; Berild and Fuglstad,, 2023) and stochastic advection-diffusion models (Clarotto et al.,, 2024) to develop more flexible models than previously considered. However, unlike Berild and Fuglstad, (2023), which uses three-dimensional space, we use two-dimensional space as computations would be too expensive with three-dimensional space and time unless spatial resolution were low.
@@
The behavior of the ocean is complex, and we need high-resolution deterministic numerical models that are based on the Navier-Stokes equations to describe it well. Such deterministic models are hard to update in real-time due to their computational complexity (Liu et al.,, 2022). In contrast, statistical models are highly effective for space-time interpolation and prediction, making them suitable for real-time applications. The challenge lies in sufficiently describing the complex spatio-temporal behavior. We can take advantage of the complex numerical models to estimate complex statistical surrogate models using the abundance of numerically simulated data. Many authors have employed such an approach to estimate models used in ocean monitoring with autonomous underwater vechicles (AUVs) but with some limitations and simplifications. Ge et al., (2023) estimated a purely spatial model through covariance functions for ocean salinity field in 3D and used an adaptive sampling strategy to choose locations for the AUVs to sample. Foss et al., (2022) used the output from a numerical model to estimate an advection-diffusion model with covariate-based advection and constant isotropic diffusion for monitoring mine tailings in the ocean. Berild et al., (2024) estimated a non-stationary and anisotropic spatial model through the SPDE approach for a 3D ocean salinity field.
In Section 2, we outline the theoretical foundation for spatio-temporal modeling using SPDEs. Then in Section 3, we introduce the proposed method for modeling spatio-temporal data and elaborate on the parameterization of the stochastic advection-diffusion equation. In Section 4, we demonstrate the proposed method on a synthetic dataset with a complex spatio-temporal structure and compare its predictive performance to other spatio-temporal models. In Section 5, we apply our method to emulate a numerical ocean model, simulate an ocean monitoring scenario with AUVs, and compare the accuracy of the proposed model to a non-stationary separable spatio-temporal model. Finally, Section 6 presents a discussion of the results and our conclusions. Some technical details such as the derivation of the numerical solution of the stochastic advection-diffusion equation are deferred to the appendices. The computer code used in this work is available at https://github.com/berild/spdepy.
@@ 2.1 Non-stationary spatial modelling
where $\kappa(\cdot)$ is a positive function, $\nabla$ is the gradient, and $\mathbf{H}(\cdot)$ is a differentiable spatially varying symmetric positive definite $2\times 2$ matrix, and $\mathcal{W}(\cdot)$ is Gaussian white noise. Note that $\mathcal{W}(\cdot)$ is a slight abuse of notation as Gaussian white noise must be viewed as a generalized Gaussian random field that does not make sense point-wise; see, e.g., Adler and Taylor, (2007, Section 1.4.3). See Bolin and Kirchner, (2020) for more details on the technical requirements on $\kappa(\cdot)$ and $\mathbf{H}(\cdot)$. We apply zero-flow Neumann boundary conditions,
@@
In the case that $\kappa(\cdot)$ and $\mathbf{H}(\cdot)$ are spatially varying, there is no obvious way to avoid boundary effects since there is no unique way to extend $\kappa(\cdot)$ and $\mathbf{H}(\cdot)$ to a larger domain than the area of interest. In practice, an option is to use a larger domain than the area of interest, and then estimate $\kappa(\cdot)$ and $\mathbf{H}(\cdot)$ also outside the area of interest (Fuglstad et al., 2015b, ). The anisotropic Laplacian in Equation (1) induces local anisotropy in the correlation at every location, and the combination of $\kappa(\cdot)$ and $\mathbf{H}(\cdot)$ controls the spatially varying spatial dependence and spatially varying marginal variance. See the discussion in Fuglstad et al., 2015b .
In this paper, we will follow Fuglstad et al., 2015a and parametrize $2\times 2$ symmetric positive-definite matrices through
@@ 2.2 Separable spatio-temporal modelling
$\frac{\partial}{\partial t}u(\bm{s},t)+cu(\bm{s},t)=\sqrt{2c}\frac{\partial}{\partial t}B(\bm{s},t),\quad\bm{s}\in\mathcal{D},\quad t\in(0,\infty),$ (5)
@@ 2.3 The stochastic advection-diffusion equation
$\left(-\mathbf{H}_{\mathrm{E}}(\bm{s})\nabla u(\bm{s},t)+\bm{\omega}_{\mathrm{E}}(\bm{s})u(\bm{s},t)\right)\cdot{\bm{n}}(\bm{s})=0,\quad\bm{s}\in\partial\mathcal{D},\quad t\in(0,\infty),$ (7)
@@ 3.2 Parametrizing the advection-diffusion-reaction model
where $B_{x,i}(x)$ and $B_{y,j}(y)$ are the 2nd-order B-spline basis functions in the $x$- and $y$-directions, respectively. A visualization of the B-spline basis functions is shown in Figure 1. This is similar to the approach taken by Fuglstad et al., 2015b and Berild and Fuglstad, (2023) on the Whittle-Matérn SPDE. This gives in total 90 coefficients for the basis functions, and thus 91 parameters including $\log(\tau)$. In general, one could use an arbitrary number of B-spline basis functions in each dimension at the cost of increased computation time. We refer readers to Berild and Fuglstad, (2023, Appendix A.3) for a detailed description of the construction of 1D clamped B-splines.
@@
In this paper, we use numerical simulations that are fully observed in space and time, and we have repeated realizations by considering numerical simulations from different time segments. For spatial non-stationarity, the work by Berild and Fuglstad, (2023) indicates that dense observations and multiple realizations are necessary when estimating so flexible models without penalties on the parameters. As demonstrated in Fuglstad et al., 2015b ; Fuglstad et al., (2019), one can add penalities or priors on the latent coefficients so that the model prefers the simpler stationary model and use sparsely observed data, but it is challenging to choose the penalization hyperparameters. Further work is needed for guidelines using flexible covariance structures in SPDE models.
We denote the parameter vector as $\bm{\theta}$, and the vector contains $\log(\tau)$ and all coefficients of basis functions. When $\bm{\theta}$ is known, all functions are known, and the stochastic advection-diffusion-reaction is fully specified. We first discretize the stochastic advection-diffusion-reaction model in space and time, then state the full hierarchical model and discuss computations.
@@ 3.3 Discretizing the advection-diffusion-reaction model
We solve the advection-diffusion-reaction SPDE given in Equation (6) numerically by first discretizing in time by backward Euler (Butcher,, 2016), and then in space using a finite volume method. This is inspired by Clarotto et al., (2024), but differ in that they used a FEM discretization in space.
@@
The spatial domain ${\cal D}$ is subdivided into $K$ regular rectangular grid cells, each represented by a centroid $\bm{s}_{k}$. We assume a piece-wise constant solution with constant value within each grid cell, as is common in standard FVM. A visualization of the spatial discretization is presented in Figure S1. We refer to Appendix A for technical details and give an overall description of the results below.
@@
${\bm{u}}^{n}=\begin{bmatrix}u^{n}(\bm{s}_{1}),\ldots,u^{n}(\bm{s}_{K})\end{bmatrix}^{\mathrm{T}},\quad n=0,\ldots,T-1.$
@@
$\left(\mathbf{D}_{\mathrm{V}}+\Delta t\left(\mathbf{D}_{\mathrm{V}}\mathbf{D}_{\kappa^{2}}-\mathbf{A}_{\mathbf{H}}+\mathbf{A}_{\bm{\omega}}\right)\right){\bm{u}}^{n+1}=\mathbf{D}_{\mathrm{V}}{\bm{u}}^{n}+\tau\sqrt{\Delta t}\mathbf{D}_{\mathrm{V}}\bm{\phi}^{n},$ (12)
@@
  • $\mathbf{D}_{\mathrm{V}}=V\mathbf{I}_{K}$: a diagonal matrix containing the areas, $V$, of the cells;
@@
  • $\mathbf{A}_{\bm{\omega}}$: uses an up-wind scheme Eymard et al., (2000, Section 20.3) for advection and details are provided in Appendix A.